DIFFERENTIAL GEOMETRY/PDE SEMINAR

Monday, May 17, 2004 Thomson 234 4:00 pm (Note: special day, time and room!)

Rigidity of asymptotically hyperbolic spaces

Gang Tian (Princeton and MIT)

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.